# From shape to fate: making bacterial swarming expansion predictable

### v1

@@ Abstract
Microbial swarming on mucosal surfaces reshapes microbial communities and influences mucosal healing and antibiotic tolerance. Yet even with time-lapse microscopy and deep learning, analyses of swarming colonies remain descriptive and cannot forecast how their fronts reorganize in time. This limitation is significant because the advancing edge determines access to nutrients, host tissue and competing microbes. We recast the expansion of Enterobacter sp. SM3 swarms as a problem of morphological forecasting, and assemble SwarmEvo, a time-lapse dataset represented as boundary-resolved segmentations. TexPol–Net, a texture- and geometry-aware segmentation model, sharpens diffuse edges and preserves fingered fronts, creating a stable substrate for dynamics. On this representation, we develop Morpher, an autoregressive forecasting network with a “Morphon” memory that links local curvature to long-range temporal dependencies. Morpher outperforms leading video-prediction models in maintaining front localization and anisotropic branching, and modest segmentation improvements yield noticeably more stable forecasts. Ablations across sequence models, inference strategies and observation ratios show that attention-based architectures with structural memory best preserve dense-finger propagation. By uniting geometry-aware segmentation with morphology-level forecasting, this framework turns swarming expansion into a predictive dynamical system, enabling quantitative interrogation and potential control of microbial collectives during mucosal repair and gut ecosystem engineering.
@@ 1 Introduction
Microbial communities form dense and continually reorganizing ecological networks within and around animal hosts [27, 7, 2, 50, 38, 36]. Their spatial organization and collective movement govern population expansion, interspecies interactions and tissue homeostasis under both healthy and diseased conditions [17, 26, 22]. Among these behaviors, collective surface motility on semi–solid substrates—classically termed swarming [1, 5]—produces continually reconfiguring fronts and patterned colony architectures that both reflect coordinated cellular behaviors and reshape local host environments [11, 49, 32]. Yet despite extensive descriptive work, the spatiotemporal evolution of such structures remains poorly characterized. In particular, forecasting the future morphology of a swarming colony—one of the most dynamic forms of bacterial collective motion—is currently out of reach. A predictive framework that links present colony architecture to its future configurations would transform morphology from a descriptive endpoint into a quantitatively interrogable variable, enabling mechanistic studies of how microbial collectives respond to environmental and physiological cues.
@@
Attempts to analyze colony growth have historically been constrained by low imaging frequency and qualitative interpretation [24, 28]. Classical computer–vision pipelines partially automated colony detection and quantification [4, 9, 37, 51], but they tend to fail when boundaries blur, textures reorganize or colonies overlap. More recent analyses of active suspensions and deformable colony–fluid interfaces have revealed turbulence–like flows and curvature–dependent edge dynamics [18, 48]. However, these approaches often presume approximate symmetry or reduce colony evolution to low–dimensional order parameters, limiting their applicability to the irregular, anisotropic and burst–like propagation characteristic of swarming fronts. In parallel, deep learning has expanded biological image analysis, yielding convolutional neural network (CNN)–based colony detectors [14, 47], temporal classifiers of motility states [31] and hybrid pipelines that combine growth–curve inference with object detection [29]. Advances in imaging, such as coherent time–lapse microscopy for early species identification [43] and engineered swarming biosensors [12], further underscore the morphological richness encoded in colony structure. Yet these developments remain essentially descriptive: they identify colonies, classify motility or extract bulk parameters, but they do not predict how a colony’s contour will evolve. Even models trained to infer motility type from single frames [23] collapse inherently dynamic processes into static labels and therefore offer limited insight into future behavior.
A central obstacle is that swarming colonies violate the appearance continuity implicitly assumed by most natural–scene video models. Fronts propagate through intermittent bursts, transient asymmetries and rapid reorganizations of multi–scale texture. Curvature modulates local velocity; protrusions emerge and retract discontinuously; textures shift without preserving pixel–level coherence. Under these conditions, extrapolating visual appearance is neither stable nor biologically grounded. Predicting future morphology instead requires reasoning over geometric and biophysical representations of the colony boundary and its local environment. Such forecasts would enable in silico perturbation experiments, real–time anticipation of responses to surface biochemistry or antimicrobial stress, and strategies to steer collective migration in applications ranging from mucosal repair to microbial therapeutics and microbiome engineering. Predictions of front position and branching patterns implicitly reflected experimental responses to mucin gradients and antimicrobial stress, indicating that the learned representations encode biologically meaningful adaptation rather than merely geometric similarity.
@@
These questions crystallize in Enterobacter sp. SM3, a representative swarming commensal from the murine gut [33, 8]. SM3 reshapes intestinal microbial organization and promotes mucosal healing, whereas mutants impaired in swarming lose these beneficial functions [11]. More broadly, swarming has been implicated in antibiotic tolerance, regulation of virulence factors and robust colonization of host–associated surfaces [21, 30], with its dynamics strongly modulated by surface biochemical cues such as mucin [32]. In inflamed intestinal environments, the advancing swarming front is the actionable interface: its future position specifies where oxygen is depleted and anaerobic niches emerge, shaping downstream microbial reorganization and recovery, as shown in Fig. 1. Anticipating this position in advance therefore provides a principled basis for spatially and temporally targeted intervention, through pharmacological or environmental modulation of the local microenvironment. Despite this conceptual clarity, translating it into predictive capability remains an open challenge. No framework currently exists for forecasting the evolving contours of swarming colonies at the resolution of individual fronts. This gap limits our ability to quantify how environmental or physiological signals shape collective migration and to connect microscopic motility programs to macroscopic patterns in a predictive manner.
Here we address this gap by formulating swarming colony expansion as a problem of morphological forecasting in a geometric state space. We first assemble the Swarming Morphogenesis Evolution (SwarmEvo) dataset, a curated collection of high–resolution time–lapse recordings of Enterobacter sp. SM3 spreading across semi–solid substrates under systematically varied conditions. Each recording is processed into boundary–resolved segmentation sequences that capture the evolving colony contour, providing both a biologically faithful account of swarming dynamics and a quantitative benchmark for evaluating whether computational models can maintain geometric fidelity over time. Building on this representation, we develop TexPol–Net, a segmentation framework that integrates texture- and geometry-aware priors to recover high–fidelity fronts even when boundaries are diffuse or heavily reorganized, thereby establishing a stable morphological state space for temporal prediction. We then introduce Morpher, a spatiotemporal model that treats forecasting as contour evolution within this state space, combining local geometric cues with long–range temporal dependencies to capture both smooth curvature modulations and rapid reorganization of protrusions.
By establishing a geometric representation of colony expansion and a forecasting model capable of operating within it, SwarmEvo, TexPol–Net and Morpher together recast swarming from a descriptive phenomenon into a tractable dynamical system with predictable trajectories. This framework enables quantitative interrogation of how microbial collectives respond to environmental and physiological cues with a temporal resolution beyond the reach of current experimental tools, and it lays an analytical foundation for integrating morphological forecasting with real-time imaging and for interpreting and steering microbial collective behavior in vivo.
@@ 2 Results
2.1 A two-stage formulation of swarming morphology
The swarming expansion of Enterobacter sp. SM3 is characterized by pronounced morphological plasticity. Under identical assay conditions, colonies may evolve through irregular, dense finger-like protrusions driven by curvature instabilities at the advancing front, or settle into near-concentric growth characterized by a stabilized rim and weak directional anisotropy. These recurring regimes define a structured spectrum of morphogenetic behavior, spanning fine-scale boundary fluctuations and colony-scale organization (Fig. 1).
Figure 1: From swarming dynamics to predictive guidance. In a mouse model of colitis, Enterobacter sp. SM3 swarms along the inflamed mucosal surface. Under pharmacological, environmental, or microbiota-based modulation, forecasting the future position of the active swarming front may enable spatially targeted and temporally coordinated interventions. Such predictive capability could be coupled with experimental strategies that locally modulate oxygen availability or microbial competition at the advancing front, with potential effects on anaerobic niche formation and downstream community composition, including Muribaculaceae expansion. Time-lapse swarming assays are converted into boundary-resolved morphological masks by TexPol–Net, with Texture-Edge and Polar–Context attention, shared prototypes and mask assembly, defining observed morphological states. Morpher models their temporal evolution with a sequence backbone and Morphon memory to predict future morphologies for quantitative evaluation. Panels I–IV depict an application-oriented biological and experimental pipeline, whereas the modeling framework follows a two-stage formulation of boundary measurement and morphology-level forecasting.
To resolve these dynamics, we adopt a two-stage formulation that separates morphological measurement from temporal prediction. In Panel III (Fig. 1), time-lapse microscopy is converted into sequences of boundary-resolved morphological masks using TexPol–Net. Rather than serving as a generic segmentation backbone, TexPol–Net acts as a measurement operator that constrains the geometric degrees of freedom of the colony front, preserving thin protrusions and internal texture while suppressing illumination-induced variability. The resulting mask sequences constitute a compact and biologically interpretable representation of colony state across time.
Panel IV (Fig. 1) treats these observed morphologies as a temporal signal and models their evolution using Morpher. Operating directly in morphology space, Morpher infers latent dynamical states from the observed window and extrapolates future colony shapes beyond it. This formulation shifts the forecasting task from image synthesis to the prediction of boundary evolution, allowing local perturbations along the front to propagate coherently over long horizons. Autoregressive conditioning further permits predicted morphologies to influence subsequent growth states, reflecting the cumulative nature of swarming expansion.
By decoupling spatial characterization from temporal dynamics, the framework limits error accumulation from appearance noise and isolates the mechanisms governing shape evolution. This structure supports systematic evaluation of forecasting models across distinct morphogenetic regimes, from highly irregular fingered fronts to stabilized radial expansion, within a unified representation that preserves biologically relevant geometric structure.
2.2 Recovering colony fronts for accurate morphological forecasting
Accurate delineation of the advancing front is the prerequisite for morphological forecasting, as it encodes curvature, anisotropy, and the local direction of expansion. We therefore evaluated how different segmentation backbones reconstruct colony fronts across representative growth regimes, and quantitatively examined how boundary accuracy evolves as matching criteria become increasingly stringent (Fig. 3).
In the dense–finger regime, the front is highly corrugated and places stringent demands on boundary fidelity (Fig. 3b). YOLOv11 is able to follow the global outline, but tends to break slender fingers into shortened or discontinuous segments, effectively smoothing high-curvature protrusions and weakening the anisotropic structure of the advancing front. SAM-based models instead emphasize the central colony mass and suppress distal protrusions, causing the boundary to evolve toward a smoother, near-circular shape. In contrast, TexPol–Net remains aligned with the annotated front along both the outer envelope and fine branches, preserving uneven finger lengths and the radial bias characteristic of early expansion; YOLOv12 is the closest competitor in this regime.

### v2

@@ Abstract
Microbial swarming on mucosal surfaces reshapes microbial communities and influences mucosal healing and antibiotic tolerance. Yet even with time-lapse microscopy and deep learning, analyses of swarming colonies remain largely descriptive and lack accurate forecasting of front reorganization. This limitation is significant because the advancing edge determines access to nutrients, host tissues, and competing microbes. The expansion of Enterobacter sp. SM3 swarms is recast as a problem of morphological forecasting, and SwarmEvo is assembled as a time-lapse dataset of boundary-resolved segmentations. TexPol–Net, a texture- and geometry-aware segmentation model, sharpens diffuse edges and preserves finger-like fronts, creating a stable substrate for dynamics. On this representation, Morpher is developed as an autoregressive forecasting network with a “Morphon” memory that links local curvature to long-range temporal dependencies. Morpher outperforms leading video-prediction models in maintaining front localization and anisotropic branching, and modest segmentation improvements yield noticeably more stable forecasts. Ablations across sequence models, inference strategies, and observation ratios show that attention-based architectures with structural memory best preserve branching propagation dynamics. By uniting geometry-aware segmentation with morphology-level forecasting, this framework turns swarming expansion into a predictive dynamical system, enabling quantitative interrogation and potential control of microbial collectives during mucosal repair and gut ecosystem engineering.
Keywords: bacterial swarming; morphological forecasting; collective behavior; spatiotemporal dynamics; deep learning; predictive modeling
Significance Statement Swarming bacteria reshape their environment through a moving front that governs access to oxygen, host tissue, and competing microbes. Yet this front has largely been described rather than predicted. We show that swarming trajectories are not organized by experimental conditions but by colony-specific dynamics, shifting prediction away from condition labels toward direct measurement of the front itself. By recasting the front as a geometric state whose evolution can be forecast, our framework links measurement fidelity to predictive stability. This establishes colony morphology as a quantitative, forward-looking variable and opens a path toward anticipating—and ultimately controlling—collective microbial behavior.
@@ 1 Introduction
Microbial communities form dense and continually reorganizing ecological networks within and around animal hosts [40, 10, 4, 69, 55, 53]. Their spatial organization and collective movement shape population expansion, interspecies interactions, and tissue homeostasis in both health and disease [23, 39, 33, 20]. Among these behaviors, collective surface motility on semi-solid substrates—classically termed swarming [1, 8, 66, 29, 3, 6]—produces continually reconfiguring fronts and colony architectures that both reflect coordinated cellular behavior and actively remodel the local host environment [14, 68, 45, 25]. Although classic studies of swarm pattern formation, cell-state differentiation, and periodic colony expansion documented rich phenomenology [50, 49, 24, 28], they stopped short of treating front evolution as a predictive problem. We still lack a framework that can predict how a swarming colony will change shape over time, a capability that would turn morphology from a visual endpoint into a quantitative variable for probing how microbial collectives respond to environmental and physiological cues.
@@
This gap persists despite substantial progress in colony analysis. Early studies were limited by sparse imaging and qualitative interpretation [35, 41]. Classical computer-vision pipelines improved colony detection and counting [7, 12, 54, 71, 2, 70], but they deteriorate when boundaries blur, textures reorganize, or colonies overlap. Physical analyses of active suspensions and deformable colony interfaces have revealed turbulence-like flows and curvature-dependent edge dynamics [26, 65], yet they often assume approximate symmetry or compress growth into low-dimensional summaries that cannot capture the irregular, anisotropic, and burst-like propagation of swarming fronts. Deep learning has broadened biological image analysis [72, 37, 19, 48] through CNN-based colony detectors [17, 64], temporal classifiers of motility states [44], and hybrid detection-growth pipelines [42]. Advances in imaging, including coherent time-lapse microscopy for early species identification [60] and engineered swarming biosensors [15], further emphasize how much biological information is encoded in colony morphology. But these approaches remain fundamentally descriptive: they detect, classify, or summarize; they do not forecast contour evolution. Even single-frame predictors of motility type [34] collapse a dynamic process into a static label.
A central reason is that swarming violates the appearance continuity assumed by most natural-scene video models. Fronts advance through intermittent bursts, transient asymmetries, and rapid multiscale reorganization. Curvature modulates local speed; protrusions emerge and retract discontinuously; texture shifts without preserving pixel-level coherence. Under these conditions, extrapolating appearance is neither stable nor biologically persuasive. What carries the future is the front itself: its geometry, branching structure, and temporal continuity. In spatially expanding microbial populations, the advancing frontier constitutes a thin, dynamically active region that governs large-scale structure while the interior remains effectively frozen [21]. A useful predictive model must therefore resolve colony-specific morphology rather than rely on coarse labels, bulk summaries, or framewise appearance alone. Forecasting future shape in this setting is not an image-generation problem but a problem of front dynamics.
@@
This challenge is especially consequential in Enterobacter sp. SM3, a representative swarming commensal from the murine gut [47, 11, 27]. SM3 reshapes intestinal microbial organization and promotes mucosal healing, whereas swarming-deficient mutants lose these beneficial effects [14]. More broadly, swarming has been linked to antibiotic tolerance, virulence regulation, and robust colonization of host-associated surfaces [32, 43, 46], and its dynamics are strongly modulated by surface biochemical cues such as mucin [45]. In inflamed intestinal environments, the advancing swarming front is the actionable interface: its future position determines where oxygen is depleted and anaerobic niches emerge, thereby shaping downstream community reorganization and recovery (Figure 1). Anticipating that front in advance would provide a principled basis for spatially and temporally targeted intervention. Yet no framework currently forecasts swarming morphology at the resolution of individual fronts, leaving a gap between microscopic motility programs and macroscopic pattern evolution.
Here we address that gap by recasting swarming colony expansion as a problem of morphological forecasting in a geometric state space. We assemble the Swarming Morphogenesis Evolution (SwarmEvo) dataset, a time-lapse resource of Enterobacter sp. SM3 expanding across systematically varied semi-solid environments. We then recover boundary-resolved colony states with TexPol–Net, designed to preserve diffuse but biologically important front structures, and forecast their evolution with Morpher, a spatiotemporal model that treats prediction as contour dynamics rather than future-image synthesis. By linking boundary measurement to long-horizon front forecasting, this framework reframes swarming as a measurable and predictable dynamical process, opening a route toward quantitative interrogation—and ultimately control—of microbial collective behavior in living environments.
@@ 2 Results
2.1 A state-based formulation of swarming morphology
Swarming expansion in Enterobacter sp. SM3 spans a reproducible morphological spectrum. Even under the same experimental conditions, colonies form two distinct morphogenetic regimes: an anisotropic branching regime (finger-like fronts) and a near-concentric regime (approximately isotropic expansion) (Figure 1). These regimes reflect structured morphogenetic states rather than frame-level visual fluctuations.
We therefore formulate swarming as a two-stage problem: first, measuring the advancing front as a geometric state; second, forecasting the evolution of that state. Time-lapse images are converted into boundary-resolved masks that preserve protrusions while suppressing appearance variability, yielding a stable representation across time. Forecasting is then posed as front evolution rather than future-image synthesis. This shift moves the problem from appearance prediction to state evolution, with morphology serving as a compact state descriptor of colony dynamics.
Figure 1: From swarming dynamics to predictive guidance. In colitis, Enterobacter sp. SM3 swarms along the inflamed mucosal surface, where the advancing front governs access to oxygen and microbial competition. Anticipating its future position enables spatially and temporally targeted intervention. Time-lapse assays are converted by TexPol–Net into boundary-resolved colony states, and Morpher forecasts their evolution. This framework links the biological problem to a two-stage formulation of boundary measurement and morphology-level prediction.
2.2 Individual variability in swarming dynamics
Before evaluating forecasting models, we asked whether swarming trajectories are cleanly organized by assay conditions. They are not. In PCA of trajectory-level perimeter and area features, colonies measured under different temperatures, humidities, and agar concentrations remain extensively overlapped rather than forming distinct clusters (Figure 2a,b). The same pattern persists at the distribution level: for both perimeter and area trajectories, within- and across-condition pairwise distances nearly coincide, with overlap coefficients remaining high (0.92–0.97) and Cliff’s $\delta$ close to zero across all three variable groupings (Figure 2c,d).
Figure 2: Large individual variability and weak separability in swarming morphology. a,b PCA of trajectory-level perimeter and area features from 81 colonies across temperature, humidity, and agar concentration. Broad overlap indicates weak separability by nominal condition. c,d Pairwise-distance distributions for perimeter and area trajectories under the same groupings. Overlap between within- and across-condition distances shows that individual variation is comparable to condition-level differences.
Swarming trajectories are not organized by nominal experimental conditions. Instead, variation is dominated by colony-specific dynamics, with within- and across-condition distances remaining statistically indistinguishable. This eliminates condition labels as a predictive axis and shifts the problem to geometry-resolved state evolution.
2.3 Segmentation defines the predictive state
Because forecasting depends on front geometry, we first evaluated how accurately different segmentation backbones reconstruct colony boundaries across representative growth regimes (Figure 3b). The clearest separation appears in the anisotropic branching regime, where boundary fidelity is most demanding. YOLOv11 captures the global outline but shortens or fragments slender fingers. SAM and SAM2 preserve the colony core while suppressing distal protrusions, driving the front toward a smoother, more circular contour. TexPol–Net remains aligned with both the outer envelope and fine branches, with YOLOv12 as the closest competitor.
Figure 3: TexPol–Net improves colony-front segmentation by coupling texture-sensitive boundary encoding with a geometry-aligned context prior. a TexPol–Net architecture within a prototype-based instance segmentation pipeline. TEA in the backbone preserves boundary texture, and PCA in the bidirectional neck maintains polar consistency during multi-scale fusion. b Qualitative comparison on representative anisotropic branching and near-concentric regimes. TexPol–Net better preserves branch heterogeneity and boundary position than YOLOv11 [30], SAM [31], and SAM2 [51], with YOLOv12 [58] as the closest competitor. c AP as a function of IoU threshold on the SwarmEvo segmentation test set (78 colonies), highlighting stronger performance under boundary-stringent matching. d Image-wise IoU distribution on the same test set, summarizing accuracy across samples. e Image-wise Dice distribution on the same test set, reflecting consistency and variability of segmentation quality.